A 65.5 kg snowboarder heads down a 17.5° hill that has a height of 70.0 m. If the hill is assumed to be frictionless and there is horizontal wind with a force of 25.3 N acting against the snowboarder, find the speed of the snowboarder as they reach the bottom of the hill using work and energy.
Answer #1

Solution) m = 65.5 kg

Angle , theeta = 17.5°

h = 70 m

Horizontal wind force , F =25.3 N

V = ?

Net force , Fnet = F' - F

F' = mg(sin(theeta))

F' = 65.5×9.8×sin(17.5°)

F' = 193.02 N

Fnet = 193.02 - 25.3 = 167.72 N

Work energy theorem

W = KE

(Fnet)(L) = (1/2)(m)(V^2)

L = h/sin(theeta)

L = (70)/sin(17.5)

L = 232.78 m

(167.72)(232.78) = (1/2)(65.5)(V^2)

V^2 = 1192.11

V = (1192.11)^(1/2)

V = 34.5 m/s